# Strategic Informed Trading and the Value of Private Information

### v1

@@ Appendix D Welfare formulas used in Section 3
$\begin{split}&\frac{1}{2}-\frac{(1-R)\widehat{y}^{2}(\beta+R(1+\widehat{y})^{2})}{2\beta(1+2\widehat{y})^{2}}-\frac{\beta+(1+\widehat{y})^{2}}{(1+2\widehat{y})(1+\widehat{y})}\\ &\qquad+\frac{\beta\lambda_{I}^{2}(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})}{2\mathbf{V}}\left(\frac{(1-R)\widehat{y}}{\beta(1+2\widehat{y})}+\frac{\lambda_{U}}{\lambda_{I}(1+\widehat{y})}\right)^{2}\\ &\qquad\qquad=\frac{\beta\lambda_{I}^{2}(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})}{2\mathbf{V}(1+\widehat{y})^{4}},\end{split}$
which if true, would finish the proof. Let us first focus on the terms with $(1+2\widehat{y})^{2}$ in the denominator on the left side. Here, using (90) we find
$\begin{split}&-\frac{(1-R)\widehat{y}^{2}(\beta+R(1+\widehat{y})^{2})}{2\beta(1+2\widehat{y})^{2}}+\frac{\beta\lambda_{I}^{2}(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})(1-R)^{2}\widehat{y}^{2}}{2\mathbf{V}\beta^{2}(1+2\widehat{y})^{2}}\\ &\qquad=-\frac{\lambda_{U}^{2}R(1-R)\widehat{y}^{2}(\beta+R(1+\widehat{y})^{2})}{2\mathbf{V}}.\end{split}$
Next, let us focus on the terms with $(1+\widehat{y})(1+2\widehat{y})$ in the denominator on the left side.
$\begin{split}&-\frac{\beta+(1+\widehat{y})^{2}}{(1+2\widehat{y})(1+\widehat{y})}+\frac{\lambda_{U}(1-R)\lambda_{I}\widehat{y}(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})}{\mathbf{V}(1+2\widehat{y})(1+\widehat{y})}\\ &\qquad=\frac{\beta+(1+\widehat{y})^{2}}{\mathbf{V}(1+2\widehat{y})(1+\widehat{y})}\bigg{(}-\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}-\lambda_{I}^{2}(1-R)(\beta+(1+\widehat{y})^{2})\\ &\qquad\qquad+\lambda_{U}\lambda_{I}(1-R)\widehat{y}(\beta+R(1+\widehat{y})^{2})\bigg{)}\end{split}$
Using (83) the terms within the parentheses evaluate to
$\begin{split}&-\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}-\lambda_{I}(1-R)(\lambda_{U}R\widehat{y}(1+\widehat{y})^{2}-\lambda_{U}\beta(1+\widehat{y}))+\lambda_{U}\lambda_{I}(1-R)\widehat{y}(\beta+R(1+\widehat{y})^{2})\\ &\qquad=\lambda_{U}\beta(1+2\widehat{y})(\lambda_{I}(1-R)-\lambda_{U}R(1+2\widehat{y})),\end{split}$
so that
$\begin{split}&-\frac{\beta+(1+\widehat{y})^{2}}{(1+2\widehat{y})(1+\widehat{y})}+\frac{\lambda_{U}(1-R)\lambda_{I}\widehat{y}(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})}{\mathbf{V}(1+2\widehat{y})(1+\widehat{y})}\\ &\qquad=\frac{\lambda_{U}\beta(\beta+(1+\widehat{y})^{2})(\lambda_{I}(1-R)-\lambda_{U}R(1+2\widehat{y}))}{\mathbf{V}(1+\widehat{y})}\end{split}$
Therefore, we must show
@@
$\begin{split}&\frac{1}{2}-\frac{\lambda_{U}^{2}R(1-R)\widehat{y}^{2}(\beta+R(1+\widehat{y})^{2})}{2\mathbf{V}}+\frac{\lambda_{U}\beta(\beta+(1+\widehat{y})^{2})(\lambda_{I}(1-R)-\lambda_{U}R(1+2\widehat{y}))}{\mathbf{V}(1+\widehat{y})}\\ &\qquad+\frac{\beta(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})}{2\mathbf{V}(1+\widehat{y})^{4}}\left(\lambda_{U}^{2}(1+\widehat{y})^{2}-\lambda_{I}^{2}\right)=0.\end{split}$
Let us focus on terms containing $\beta(\beta+(1+\widehat{y})^{2})$:
$\begin{split}&\frac{\beta(\beta+(1+\widehat{y})^{2})}{2\mathbf{V}(1+\widehat{y})^{4}}\bigg{(}2\lambda_{U}(\lambda_{I}(1-R)-\lambda_{U}R(1+\widehat{y})-\lambda_{U}R\widehat{y})(1+\widehat{y})^{3}\\ &\qquad+(\beta+R(1+\widehat{y})^{2})(\lambda_{U}^{2}(1+\widehat{y})^{2}-\lambda_{I}^{2})\bigg{)}\end{split}$
The terms within the parentheses are
@@
$\begin{split}&2\lambda_{U}\lambda_{I}(1-R)(1+\widehat{y})^{3}-2\lambda_{U}^{2}R(1+\widehat{y})^{4}-2\lambda_{U}^{2}R\widehat{y}(1+\widehat{y})^{3}+\lambda_{U}^{2}\beta(1+\widehat{y})^{2}-\beta\lambda_{I}^{2}\\ &\qquad\qquad+\lambda_{U}^{2}R(1+\widehat{y})^{4}-\lambda_{I}^{2}R(1+\widehat{y})^{2}\\ &\qquad=-\lambda_{U}^{2}R(1+\widehat{y})^{4}+2\lambda_{U}\lambda_{I}(1-R)(1+\widehat{y})^{3}-2\lambda_{U}(1+\widehat{y})(\lambda_{I}(1+\widehat{y})^{2}+\lambda_{I}\beta+\lambda_{U}\beta(1+\widehat{y}))\\ &\qquad\qquad+\lambda_{U}^{2}\beta(1+\widehat{y})^{2}-\beta\lambda_{I}^{2}+\lambda_{U}^{2}R(1+\widehat{y})^{4}-\lambda_{I}^{2}R(1+\widehat{y})^{2}\\ &\qquad=-\lambda_{U}^{2}R(1+\widehat{y})^{4}-2\lambda_{U}\lambda_{I}R(1+\widehat{y})^{3}-(\lambda_{U}^{2}\beta+\lambda_{I}^{2}R)(1+\widehat{y})^{2}\\ &\qquad\qquad-2\lambda_{U}\lambda_{I}\beta(1+\widehat{y})-\beta\lambda_{I}^{2}\\ &\qquad=-(\beta+R(1+\widehat{y})^{2}\left(\lambda_{U}^{2}(1+\widehat{y})^{2}+2\lambda_{U}(1+\widehat{y})\lambda_{I}+\lambda_{I}^{2}\right)\\ &\qquad=-(\beta+R(1+\widehat{y})^{2})(1+\lambda_{U}\widehat{y})^{2}.\end{split}$
Above, the second equality uses (83). Therefore, the terms containing $\beta(\beta+(1+\widehat{y})^{2})$ are
$-\frac{\beta(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})(1+\lambda_{U}\widehat{y})^{2}}{2\mathbf{V}(1+\widehat{y})^{4}},$
and we must show
@@
$\begin{split}0&=\frac{1}{2}-\frac{\lambda_{U}^{2}R(1-R)\widehat{y}^{2}(\beta+R(1+\widehat{y})^{2})}{2\mathbf{V}}-\frac{\beta(\beta+(1+\widehat{y})^{2})(\beta+R(1+\widehat{y})^{2})(1+\lambda_{U}\widehat{y})^{2}}{2\mathbf{V}(1+\widehat{y})^{4}},\end{split}$
or that
$\begin{split}\mathbf{V}&=(\beta+R(1+\widehat{y})^{2})\bigg{(}\lambda_{U}^{2}R(1-R)\widehat{y}^{2}+\frac{\beta(\beta+(1+\widehat{y})^{2})(1+\lambda_{U}\widehat{y})^{2}}{(1+\widehat{y})^{4}}\bigg{)}.\end{split}$
To show this, we note that (69) implies
$\beta=\frac{(1+\widehat{y})^{2}(\lambda_{U}R\widehat{y}-\lambda_{I})}{1+\lambda_{U}\widehat{y}},$
and hence from (90) we see
@@
$\mathbf{V}=(1+\widehat{y})^{2}\left(\lambda_{I}^{2}(1-R)+\frac{(\lambda_{U}R\widehat{y}-\lambda_{I})(\lambda_{U}^{2}R(1+2\widehat{y})^{2}+\lambda_{I}^{2}(1-R))}{1+\lambda_{U}y}\right),$
and
@@
$\begin{split}&(\beta+R(1+\widehat{y})^{2})\bigg{(}\lambda_{U}^{2}R(1-R)\widehat{y}^{2}+\frac{\beta(\beta+(1+\widehat{y})^{2})(1+\lambda_{U}\widehat{y})^{2}}{(1+\widehat{y})^{4}}\bigg{)}\\ &\qquad=(1+\widehat{y})^{2}\frac{\lambda_{U}R\widehat{y}-\lambda_{I}+R(1+\lambda_{U}\widehat{y})}{1+\lambda_{U}\widehat{y}}\bigg{(}\lambda_{U}^{2}R(1-R)\widehat{y}^{2}+(\lambda_{U}R\widehat{y}-\lambda_{I})^{2}+\\ &\qquad\qquad+(\lambda_{U}R\widehat{y}-\lambda_{I})(1+\lambda_{U}\widehat{y})\bigg{)}.\end{split}$
Thus, we want to show
$\begin{split}&\lambda_{I}^{2}(1-R)(1+\lambda_{U}y)+(\lambda_{U}R\widehat{y}-\lambda_{I})(\lambda_{U}^{2}R(1+2\widehat{y})^{2}+\lambda_{I}^{2}(1-R))\\ &\qquad=(\lambda_{U}R\widehat{y}-\lambda_{I}+R(1+\lambda_{U}\widehat{y}))\bigg{(}\lambda_{U}^{2}R(1-R)\widehat{y}^{2}+(\lambda_{U}R\widehat{y}-\lambda_{I})^{2}+\\ &\qquad\qquad+(\lambda_{U}R\widehat{y}-\lambda_{I})(1+\lambda_{U}\widehat{y})\bigg{)}\\ &\qquad=(2\lambda_{U}R\widehat{y}+R-\lambda_{I})\bigg{(}2\lambda_{U}^{2}R\widehat{y}^{2}+(-2\lambda_{U}R\lambda_{I}+\lambda_{U}R-\lambda_{U}\lambda_{I})\widehat{y}+\lambda_{I}^{2}-\lambda_{I}\bigg{)}\\ \end{split}$
Here, as we no longer need appeal to (69) or (83) we simply match powers of $\widehat{y}$ on both sides of the equation. The cubic terms are
$4\lambda_{U}^{3}R^{2}\textrm{ vs. }4\lambda_{U}^{3}R^{2}\qquad\checkmark.$
The quadratic terms are
$\begin{split}&4\lambda_{U}^{3}R^{2}-4\lambda_{U}^{2}\lambda_{I}R\textrm{ vs. }2\lambda_{U}^{2}R(R-\lambda_{I})+2\lambda_{U}R(-2\lambda_{U}R\lambda_{I})+\lambda_{U}R-\lambda_{U}\lambda_{I})\qquad\checkmark.\end{split}$
The linear terms are
$\begin{split}&\lambda_{U}\lambda_{I}^{2}(1-R)+\lambda_{U}R(\lambda_{I}^{2}(1-R)+\lambda_{I}^{2}R)-4\lambda_{I}\lambda_{U}^{2}R\textrm{ vs. }\\ &\qquad 2\lambda_{U}R(\lambda_{I}^{2}-\lambda_{I})+(R-\lambda_{I})(-2\lambda_{U}R\lambda_{I}+\lambda_{U}R-\lambda_{U}\lambda_{I})\end{split}$
Within this term, the quadratic powers in $R$ are
$\begin{split}-\lambda_{U}\lambda_{I}^{2}+\lambda_{U}^{3}\textrm{ vs. }-2\lambda_{U}\lambda_{I}+\lambda_{U}\qquad\checkmark.\end{split}$
The linear powers in $R$ are
$-4\lambda_{U}^{2}\lambda_{I}\textrm{ vs. }4\lambda_{U}\lambda_{I}^{2}-4\lambda_{U}\lambda_{I}\qquad\checkmark.$
The constants are
$\lambda_{U}\lambda_{I}^{2}\textrm{ vs. }\lambda_{U}\lambda_{I}^{2}\qquad\checkmark.$
Lastly, terms which do not depend on $\widehat{y}$ are
$\begin{split}&\lambda_{I}^{2}(1-R)-\lambda_{I}(\lambda_{U}^{2}R+\lambda_{I}^{2}(1-R))=-\lambda_{U}\lambda_{I}R+\lambda_{U}\lambda_{I}^{2}\textrm{ vs.}\\ &\qquad R(\lambda_{I}^{2}-\lambda_{I})-\lambda_{I}^{2}(\lambda_{I}-1)=-\lambda_{U}\lambda_{I}R+\lambda_{U}\lambda_{I}^{2}\qquad\checkmark.\end{split}$
The proof is complete. ∎
Appendix E Proofs from Section 3
Proof of Proposition 3.2.
Set $c=(1-\lambda)/\lambda$, $p=p_{I}$ and write $\widehat{y}(p_{I})=y(p)$ so that (18) is

### v2

@@ B.4. Utility formulas used in Section 3
$\begin{split}e^{-\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0-,\iota}}=\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}p_{\iota}-\widehat{\psi}_{U,\iota}^{\prime}(X-p_{\iota})}\right]&=\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}X}\right]\mathbb{E}^{{\mathbb{Q}}_{0}}\left[\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{\widehat{\theta}_{U,\iota}^{\prime}(X_{0}-p_{0,\iota})}\big{|}\sigma(H_{0,\iota})\right]\right],\\ &=e^{-\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{nsn}}\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\frac{\lambda_{I}^{2}(1-R)^{2}(\beta+(1+\widehat{y})^{2})}{2\lambda_{U}^{2}R(\beta+R(1+\widehat{y})^{2})(1+2\widehat{y})^{2}}\left|P_{X}^{1/2}H_{0,\iota}\right|^{2}}\right].\end{split}$
Using (81) and (83) we deduce
$P_{X}^{1/2}H_{0,\iota}=\sqrt{\frac{\beta+R(1+\widehat{y})^{2}}{\beta(1-R)}}\mathcal{E},$
where $\mathcal{E}\sim N(0,1_{d})$ under ${\mathbb{Q}}_{0}$. This gives
@@
$\begin{split}\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}p_{\iota}-\widehat{\psi}_{U,\iota}^{\prime}(X-p_{\iota})}\right]&=e^{-\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{nsn}}\left(1+\frac{\lambda_{I}^{2}(1-R)(\beta+(1+\widehat{y})^{2}}{\lambda_{U}^{2}\beta R(1+2\widehat{y})^{2}}\right)^{-1/2},\end{split}$
and hence the ex-ante certainty equivalent for $U$ in (90).
Price-taking certainty equivalents
Here, we consider the price-taking case, with the goal of proving the following.
Proposition B.2.
Using the notation of (77), the interim price taking certainty equivalents of Section 3 are
@@
(91) $\begin{split}\textrm{CE}^{I}_{0}(G,Z_{N})&=\textrm{CE}^{o}_{I,0}(G)\\ &\qquad+\frac{\alpha_{I}R}{2(\lambda_{I}+\beta+R\lambda_{U})^{2}}\left|(1-R)\lambda_{U}P_{X}^{1/2}(G-p_{{\mathbb{Q}}_{0}})-(\lambda_{I}+\beta)P_{X}^{-1/2}\frac{Z_{N}}{\alpha_{I}}\right|^{2},\\ \textrm{CE}^{U}_{0}(H)&=\textrm{CE}^{o}_{U,0}(H_{\iota})+\frac{\alpha_{U}\lambda_{I}^{2}(1-R)^{2}R(1+\beta)}{2(\lambda_{I}+\beta+R\lambda_{U})^{2}(R+\beta)}\left|P_{X}^{1/2}(H-p_{{\mathbb{Q}}_{0}})\right|^{2}.\end{split}$
The ex-ante price taking certainty equivalents of Section 3 are
@@
(92) $\begin{split}\textrm{CE}^{I}_{0-}&=\textrm{CE}^{I}_{nsn}+\frac{\alpha_{I}}{2}\log\left(1+\frac{(1-R)(\lambda_{U}^{2}\beta R+(\lambda_{I}+\beta)^{2})}{\beta(\lambda_{I}+\beta+R\lambda_{U})^{2}}\right),\\ \textrm{CE}^{U}_{0-}&=\textrm{CE}^{U}_{nsn}+\frac{\alpha_{U}}{2}\left(1+\frac{\lambda_{I}^{2}(1-R)R(1+\beta)}{\beta(\lambda_{I}+\beta+R\lambda_{U})^{2}}\right).\end{split}$
Proof of Proposition B.2.
We start with the insider. From (43), (47) and (64)
$\begin{split}e^{-\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0}}&=\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}X}\big{|}\sigma(G,Z_{N})\right]\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{\widehat{\theta}_{I}^{\prime}(X_{0}-p_{0})}\big{|}\sigma(G_{0},Z_{N})\right].\end{split}$
From (65), (76), (77) and (84) we obtain after some simplification
@@
$\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{\widehat{\theta}_{I}^{\prime}(X_{0}-p_{0})}\big{|}\sigma(G_{0},Z_{N})\right]=e^{-\frac{R}{(\lambda_{I}+\beta+R\lambda_{U})^{2}}\left|(1-R)\lambda_{U}P_{X}^{1/2}G_{0}-(\lambda_{I}+\beta)P_{X}^{-1/2}\frac{Z_{N}}{\alpha_{I}}\right|^{2}}.$
Therefore, using (87) we obtain the interim certainty equivalent for $I$ in (91). For the ex-ante certainty equivalent we must compute
$e^{-\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0-}}=\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}p-\widehat{\psi}_{I}^{\prime}(X-p)}\right]=\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}X}\right]\mathbb{E}^{{\mathbb{Q}}_{0}}\left[\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{\widehat{\theta}_{I}^{\prime}(X_{0}-p_{0})}\big{|}\sigma(G_{0},Z_{N})\right]\right].$
Using (88) and the above we obtain
$e^{-\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0-}}=e^{-\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{nsn}}\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\frac{R}{(\lambda_{I}+\beta+R\lambda_{U})^{2}}\left|(1-R)\lambda_{U}P_{X}^{1/2}G_{0}-(\lambda_{I}+\beta)P_{X}^{-1/2}\frac{Z_{N}}{\alpha_{I}}\right|^{2}}\right].$
If we re-write this using the notation of (77), (81) we obtain
$e^{-\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0-}}=e^{-\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{nsn}}\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\frac{R}{(\lambda_{I}+\beta+R\lambda_{U})^{2}}\left|(1-R)\lambda_{U}\left(\mathcal{E}_{X}+\sqrt{\frac{R}{1-R}}\mathcal{E}_{I}\right)-(\lambda_{I}+\beta)\sqrt{\frac{1-R}{\beta R}}\mathcal{E}_{N}\right|^{2}}\right].$
We can write
$(1-R)\lambda_{U}\left(\mathcal{E}_{X}+\sqrt{\frac{R}{1-R}}\mathcal{E}_{I}\right)-(\lambda_{I}+\beta)\sqrt{\frac{1-R}{\beta R}}\mathcal{E}_{N}=\sqrt{\frac{(1-R)(\lambda_{U}^{2}\beta R+(\lambda_{I}+\beta)^{2})}{\beta R}}\mathcal{E},$
where $\mathcal{E}\sim N(0,1_{d})$ under ${\mathbb{Q}}_{0}$ $N(0,1_{d})$. This implies
$e^{-\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{0-}}=e^{-\frac{1}{\alpha_{I}}\textrm{CE}^{I}_{nsn}}\left(1+\frac{(1-R)(\lambda_{U}^{2}\beta R+(\lambda_{I}+\beta)^{2})}{\beta(\lambda_{I}+\beta+R\lambda_{U})^{2}}\right)^{-1/2},$
which gives the formula in (92). We now consider the uninformed trader. From (43), (47) and (64)
$e^{-\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0}}=\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}X}\big{|}\sigma(H)\right]\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{\widehat{\theta}_{U}^{\prime}(X_{0}-p_{0})}\big{|}\sigma(H_{0})\right].$
From (76) (with $P_{U}$ replacing $P_{I}$ and $H_{0}$ replacing $G_{0}$), (77), (79) and (84) we obtain after some simplification
$\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{\widehat{\theta}_{U}^{\prime}(X_{0}-p_{0})}\big{|}\sigma(H_{0})\right]=e^{-\frac{\lambda_{I}^{2}(1-R)^{2}R(1+\beta)}{2(\lambda_{I}+\beta+R\lambda_{U})^{2}(R+\beta)}\left|P_{X}^{1/2}H_{0}\right|^{2}}.$
Therefore, using (87) we obtain the interim certainty equivalent in (91). As for the ex-ante certainty equivalent, we must compute
$e^{-\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0-}}=\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}p-\widehat{\psi}_{U}^{\prime}(X-p)}\right]=\mathbb{E}\left[e^{-\widehat{\Pi}^{\prime}X}\right]\mathbb{E}^{{\mathbb{Q}}_{0}}\left[\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{\widehat{\theta}_{U}^{\prime}(X_{0}-p_{0})}\big{|}\sigma(H_{0})\right]\right].$
Using (88) and the above this is
$e^{-\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0-}}=e^{-\frac{1}{\alpha_{U}}\textrm{CE}^{I}_{nsn}}\mathbb{E}^{{\mathbb{Q}}_{0}}\left[e^{-\frac{\lambda_{I}^{2}(1-R)^{2}R(1+\beta)}{2(\lambda_{I}+\beta+R\lambda_{U})^{2}(R+\beta)}\left|P_{X}^{1/2}H_{0}\right|^{2}}\right].$
From (81), (82) we can write
$H_{0}=\sqrt{\frac{\beta+R}{\beta(1-R)}}P_{X}^{-1}\mathcal{E},$
where $\mathcal{E}\sim N(0,1_{d})$ under ${\mathbb{Q}}_{0}$. This gives
@@
$\begin{split}e^{-\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{0-}}=e^{-\frac{1}{\alpha_{U}}\textrm{CE}^{U}_{nsn}}\left(1+\frac{\lambda_{I}^{2}(1-R)R(1+\beta)}{\beta(\lambda_{I}+\beta+R\lambda_{U})^{2}}\right)^{-1/2},\end{split}$
and hence the ex-ante certainty equivalent for $U$ in (92). ∎
Remark B.3.
If we move from the notation of (77) to the notation in (7) we obtain the ex-ante certainty equivalents
$\begin{split}\textrm{CE}^{I}_{0-,\iota}&=\textrm{CE}_{nsn}^{I}+\frac{\alpha_{I}}{2}\log\left(1+\frac{\kappa p_{I}(1+p_{I})+\widehat{y}^{2}}{\kappa(1+p_{I})(1+2\widehat{y})}\right),\\ \textrm{CE}^{U}_{0-,\iota}&=\textrm{CE}_{nsn}^{U}+\frac{\alpha_{U}}{2}\log\left(1+\frac{\lambda^{2}(\kappa p_{I}+(1+\widehat{y})^{2})}{(1-\lambda)^{2}\kappa(1+2\widehat{y})^{2}}\right),\\ \textrm{CE}^{I}_{0-}&=\textrm{CE}_{nsn}^{I}+\frac{\alpha_{I}}{2}\log\left(1+\frac{(1-\lambda)^{2}\kappa p+(1+p_{I})(\lambda+\kappa p_{I})^{2}}{\kappa(1+\lambda p_{I}+\kappa p_{I}(1+p_{I}))^{2}}\right),\\ \textrm{CE}^{U}_{0-}&=\textrm{CE}_{nsn}^{U}+\frac{\alpha_{U}}{2}\log\left(1+\frac{\lambda^{2}(1+\kappa p_{I})}{\kappa(1+\lambda p_{I}+\kappa p_{I}(1+p_{I}))^{2}}\right).\end{split}$
B.5. Proofs from Section 3
Proof of Theorem 3.2.
Set $c=(1-\lambda)/\lambda$, $p=p_{I}$, $\widehat{y}(p_{I})=y(p)$ so (6) is